Find the limit of 2x(x-2)/x-2 as x is 2

1
Expert's answer
2021-11-29T04:57:02-0500

"lim_{x\to2}\frac{2x(x-2)}{(x-2)}\space[0/0 form]\\
=lim_{x\to2}\frac{2x+2(x-2)}{1}\\
=lim_{x\to2}\frac{4x-4}{1}\\
=4"